# Water-rich sub-Neptunes and rocky super Earths around different Stars: Radii shaped by Volatile Partitioning, Formation, and Evolution

### v1

@@ 2.2 Evolution
After the planets have accreted solids and gas and might have migrated, the protoplanetary disk eventually dissipates. Starting from this point in time, defined by the local disk pressure around the planet dropping to low levels, we model the evolution of the planet as individual body. In this work, we modified the approach of Emsenhuber et al. (2021) in a few key aspects similar to Burn et al. (2024) in order to better model planets in the sub-Neptune regime with large volatile contents. The main aspect is the used equation of state (EOS) for volatiles for which we use an EOS for water (Haldemann et al., 2020) which covers all phases. For simplicity, we treat for the evolution phase all volatile species accreted as ice (i.e. H2O, CO, CO2, CH3OH, CH4, NH3, N2, and H2S, see Marboeuf et al., 2014a, b) as water ice and will use ”water” and ”volatiles” as synonyms in this work.
@@ 2.2.1 Planetary structure and composition
In all cases, we start at the disk dissipation time with initial conditions from the formation part of the models (Sect. 2.1). To improve the consistency, we recover the luminosities of the planets at analytically and rerun the evolution part from this time onward. A difference to Burn et al. (2024) is that we also use the hydrogen and helium (H/He) mass from as starting condition instead of the one after of evolution. This has an effect on simulated planets for which an impact occurred after . By starting with , we assume that no H/He is lost for the more massive planet in a collision while all H/He is lost for the impacting planet. The mass of other components, that is iron, rocky materials, and volatile ices, is however assumed to perfectly merge. Future works will re-introduce a more realistic impact stripping with the possibility of mass loss of all species. Therefore, in the scenarios where we assume that volatile species mix with H/He (Mixed, Fractionation, see below) we obtain a total heavy element content (by mass) in the envelope of
smaller or equal to that in Burn et al. (2024).
Further common elements in all models are the presence of an iron core with a silicate MgSiO3 (perovskite) mantle both materials described with a modified polytropic equation of state of Seager et al. (2007).
In all model cases, the uppermost – third or fourth – layer is what we call the envelope and is numerically resolved in one dimension. Similar to the formation stage (described in Emsenhuber et al., 2021), the following internal structure equations (e.g. Kippenhahn et al., 2012) in hydrostatic equilibrium are solved:
@@
is smaller than the radiative gradient. The latter depends on the local optical depth . If , we use the atmosphere solution of Guillot (2010) which considers an outgoing infrared flux and a well separated in-going flux in the visible wavelength regime in a in-parts gray atmosphere under the Eddington approximation (see also Jin et al., 2014). In case of large optical depths, radiative diffusion is a well justified approximation, thus
@@
Following Mordasini et al. (2012b), we assume that the luminosity is constant throughout the structure which reduces the number of equations and is justified as only the upper, radiative regime depends directly on the luminosity. The system of equations is closed by the EOS (see below). A shooting method is used to solve the boundary value problem which arises due to core size, mass, and temperature being given as lower boundary while external pressure and density are upper boundaries to the envelope structure. A dependency on the stellar type enters directly into the upper boundary temperature of the planetary structures in the form of the irradiation temperature , where the stellar temperature and radius are taken from the tracks of Baraffe et al. (2015).
Assuming that a hydrostatic equilibrium is established on short timescales, the structure of the planet is recalculated after each step forward in time. From the obtained pressure and temperature profile, the corresponding change in energy content (gravitational and internal) can then be estimated. This estimate can be used as the luminosity for the next timestep and leads to a method which was found to result in a consistent treatment of a cooling and contracting planet (Mordasini et al., 2012b). In addition to the energy budget from solid and gas accretion, we consider radiogenic heating from the decay of 40K, 232Th, and 238U (Mordasini et al., 2012a) as an additional energy source. We further add an empirically determined source of energy causing the observed bloating of close-in planets following Sarkis et al. (2021).
@@
As for the solid part of the planetary structure, we require an equation of state for the envelope matter. For the mixed model, the envelope is assumed to consist of a perfect mixture of water, hydrogen, and helium. Therefore, the envelope metallicity is set to . While perfect mixing at all conditions is an extreme assumption, there are clear indications that H/He and water are well miscible at the high temperatures within the runaway greenhouse limit (Vazan et al., 2022; Pierrehumbert, 2023), which is a safe assumption for the observed population of small transiting exoplanets. The equation of state for water follows Haldemann et al. (2020) while that for H/He follows Chabrier et al. (2019). To combine them, we assume that Amagat’s law of additive volumes holds.
@@
In the Layered model, water and H/He are assumed to not mix at all to explore the other fringe case. In this case, the planetary structure follows a pure H/He envelope on top of the solid ice or rock layers. The ice layer density and thermal coefficients are modeled following Seager et al. (2007). This is the case as long as H/He is present. After it is lost, we switch for the volatiles to a pure water envelope following the AQUA equation of state of Haldemann et al. (2020). Because AQUA also includes gaseous and supercritical phases of water, this transition implies for hot planets that we first force the ice to be solid, motivated by the pressure of the H/He on top of it, and then transform it to a typically supercritical vapor envelope. In our static model, this transition occurs instantaneously. Especially at short orbital periods, this introduces an inconsistency as the water would evaporate somewhat earlier even if the assumption of non-mixing holds. However, this transition is relatively short and we explore this layered case as a useful extreme assumption which complements the mixing one. We note that for planets which do not contain volatile ices or for planets without any H/He after formation the Layered and Mixed models are equivalent and that the assumption of the Layered model are the same as what is used during the early formation stage of the model.
@@
The structure and initialization of the Fractionation model is equal to the Mixed model explained above. The difference lies in loosening the assumption that the gas is perfectly mixed also in the photoevaporative wind. Instead, an analytic prescription (Hunten et al., 1987; Zahnle & Kasting, 1986; Zahnle et al., 1990) detailed in Appendix A (optimistic estimate)222We use the analytic estimates from Appendix A and the expressions for there instead of the output from the tables of Johnstone (2020) (Sect. 2.2.2) for its broader applicability. is used to determine the metallicity in the photoevaporative mass loss . Consequently, this also implies that can evolve, that is, increase, over time.
@@
A novelty in evolutionary calculations is the here presented Water Sequestration model. Instead of mixing the accreted volatile species (treated as water) mass only within the envelope, we maximize the amount of water sequestered in the deep interior by assuming a molten magma ocean. In that case, some amount of volatiles can dissolve into the magma and can even dissolve into the metallic core based on the calculations in Luo et al. (2024) (see also Dorn & Lichtenberg, 2021). The envelope volatile mass at the start of our long-term evolution phase is then only and the envelope metallicity . The rationale of the model is to provide an upper limit for the mass of volatile species, treated as water, which could be removed from the envelope mass budget. This has implications for the evolution of the planetary envelope.
The partitioning of water between the mantle and surface water layers is modeled using a modified version of Henry’s law which relates the water mass fraction in the magma to the pressure at the bottom of the gaseous envelope (see Dorn & Lichtenberg (2021) for details).
The parameters and were fitted to experimental data by Dorn & Lichtenberg (2021). Partitioning from the magma ocean to the metallic core (to determine ) follows Luo et al. (2024) who assume iron silicate equilibration at mid-mantle pressures. Their partitioning coefficients depend on the concentration of water in the molten silicates as well as mid-mantle pressures but only marginally on the temperature at a given pressure.
For the coupling to the evolutionary cooling and photoevaporation model, tabulated results from these partitioning and solubility laws and further using the interior structure model detailed in Appendix B were used. We note that we did not include an explicit temperature dependency on the water partitioning leading to a weak effect of external temperature via the pressure dependencies. Therefore, the tables were using an exterior temperature of 3000 K for all models and varying total water mass budgets (from 1% to 30%) and planetary masses (0.5 to 30 M⊕) were considered. For the equilibration of water between silicates and iron, we use mid-mantle pressures and the iron mass fraction of the dry core was assumed to be 0.244 consistent with the outcome of the equilibrium condensation model used in our formation model (Thiabaud et al., 2014). To construct the tables, the presence of H/He, which would increase pressures and thus solubilities, was excluded for simplicity but supported by the fact that most planets in the considered size regime contain little H/He in mass.
@@
During time evolution, the total mass of the planet and the atmospheric pressure can decrease due to photoevaporative mass loss. Therefore, the balance between sequestered water in the magma ocean and envelope (Eq. 7) is re-evaluated after each timestep and is re-determined using the same tables as initially. The water partitioning between envelope and magma ocean is therefore dynamically adapted to the new conditions which leads to outgassing of water from the magma. In contrast, we assume that the volatiles trapped in the iron core remain constant over time (similar to outgassed scenarios in Rogers et al. (2024)).
For better comparison to the other structure models, the Water Sequestration model calculates the average density of the volatile-enriched metallic core and magma ocean using the equation of state of Seager et al. (2007) assuming the same separation into iron core, perovskite mantle, and water ice layer as in the Layered model. This implies that for the radius and bulk density of the interior, we do not assume that sequestration to the mantle and core occurred. Instead, the sequestered water (of mass ) is treated as a high-pressure ice layer on top of the rocky material. A comparison to densities calculated consistently with the structure assumption of the Water Sequestration model (Fig. 1) is presented in Appendix B.
@@ 2.2.2 Photoevaporative mass loss
To model the escape of particles from the upper atmosphere due to high-energy radiation, we use a similar prescription for photoevaporation as in Affolter et al. (2023); Burn et al. (2024). The approach is equivalent between models where the envelope metallicity varies. We assume that the escape is in the thermally-driven hydrodynamic regime (see Gronoff et al., 2020, for a recent review). The total mass loss is a combination of the loss of H/He (Kubyshkina & Fossati, 2021) and that of volatile species treated as water (Johnstone, 2020). Given the , differing between models (see Sect. 2.2.1), it is given by
where we use the tabulated values of Kubyshkina & Fossati (2021) for . The results are based on the chemical-hydrodynamic models of Kubyshkina et al. (2018), assume a 15% heating efficiency of the high-energy radiation, and depend on stellar mass, X-ray and extreme ultraviolet (EUV) fluxes , planetary masses and radii, as well as equilibrium temperature. To transition from these models applicable for pure H/He planets to water-dominated cases, the results of Johnstone (2020) are used to determine in Eq. (8) by adjusting the efficiency parameter in the energy-limited mass loss equation
where is the received flux at the planet location in either X-ray (dominating early) or EUV wavelengths, is the radius of the layer where an optical depth of 2/3 is reached from the inside-out, and is the base of the ionization layer. It is found by searching within the resolved planetary structure for the location where an optical depth of one if reached for UV photons (Murray-Clay et al., 2009). The factor , with being the ratio of the planets Roche limit to its radius. As in Burn et al. (2024), the time evolution of follows (McDonald et al., 2019) and we linearly interpolate or extrapolate to lower stellar masses if required. In contrast to Burn et al. (2024), we reduced the rates of Johnstone (2020) to account for an additional cooling effect fitted following Yoshida et al. (2022) as , where is the unmodified rate fit to the results of Johnstone (2020).
@@ 2.3 Observational data and sampling
To compare simulations and observations, we use the collection of well characterized exoplanets from Parc et al. (2024) with the goal of comparing mass-radius relations. For our purposes, we restrict ourselves to planets within 30 days orbital periods and smaller than 7 R⊕. Furthermore we require precisely determined masses and radii with relative errors below 25% and 8% respectively. This results in a set of 201 observed planets. As a second exploration, we will relate the Solar-mass population of simulated exoplanet radii to the data from the California-Kepler survey (Fulton & Petigura, 2018). For this second set of data, no precise masses are available but the application of the bias from Kepler can be done in a rigorous fashion using the KOBE package Mishra et al. (2021).
@@
To allow for a tentative, quantitative comparison also for the sample from Parc et al. (2024), we apply an estimate for the observational bias and re-sample the synthetic population that is an outcome of the model described above such that the distribution of stellar masses and the right number of observed stars is reproduced. The observational data originates from heterogeneous sources and a human intervention bias is introduced when follow-up characterization observations are scheduled. Nevertheless, a significant fraction of the planets were discovered by TESS (Ricker et al., 2014) which is a transit survey. Therefore, we calculate the geometrical transit probability (Petigura et al., 2018)
@@
and an estimate for the probability that a transiting planet is also detected (based on the Kepler mission, as detailed by Fulton et al., 2017; Petigura et al., 2018)
@@
where is the standard cumulative distribution function, , and . Here, we use the expected scaling of signal to noise with observation time (Petigura et al., 2018), estimate the mission duration ratio of Kepler to TESS , and increase by a factor of two to account for the heterogeneous sample as well as the variability in TESS observation duration per start (depends on the sector). We visually verify the estimated detection probabilities by comparing them to the results of the detailed analysis of TESS planet detection yields by Kunimoto et al. (2022).
@@
Most of the models (except the Layered one) rely on the assumption that water mixes with H/He. This is established for planets which are interior to the runaway greenhouse limit where water cannot condense due to too high temperatures. Therefore, we apply a cut at the runaway greenhouse limit which affects planets around different stars differently: The runaway greenhouse limit lies at close orbits similar to the detection thresholds for late M dwarfs only. For the estimate we use either the observed luminosities for the observational data or for synthetic stars, we use stellar luminosities at from the tracks by Baraffe et al. (2015). The incident flux of shortwave radiation on the planetary surface is on average , where is the planets’ albedo. Assuming that the planet is in thermal equilibrium and for a limiting case assuming no internal heat content, we can equate this to a critical outgoing long-wave radiation limit of from Boukrouche et al. (2021). The resulting critical distance
@@
is then a conservative upper limit within which our assumption that the water evaporates and mixes with other constituents holds. To remain conservative, we use a relatively high albedo of motivated by the results of Kopparapu et al. (2014) for high-pressure envelopes.
@@
To generate a synthetic sample of planets we sample systems from the original synthetic data weighted by stellar mass chosen such that the stellar mass distribution of the synthetic data matches the observed one (equivalent to Schlecker et al., 2022). We then proceed to draw “observed” planets by repeatedly (1000 times a set of 201) sampling planets weighted by their individual thus allowing for an individual, modeled planet to be present several times in the synthetic comparison sample. We note that this approach ignores possible correlations of the detection probability within a planetary system and we used discrete stellar mass bins (0.1, 0.3, 0.5, 0.7, 1.0, and ) when the synthetic data was created (Burn et al., 2021).
@@
Since the observed planets, to which we want to compare our models, are also characterized by radial velocity measurements, we further require for planets to be included in the synthetic comparison data that the semi-amplitude (e.g. Cumming et al., 1999, their Eq. 1) is larger than .

### v2

@@ 2.2 Evolution
After the planets have accreted solids and gas and might have migrated, the protoplanetary disk eventually dissipates. Starting from this point in time, defined by the local disk pressure around the planet dropping to low levels, we model the evolution of the planet as individual body. In this work, we modified the approach of Emsenhuber2020a in a few key aspects similar to Burn2024 in order to better model planets in the sub-Neptune regime with large volatile contents. The main aspect is the used equation of state (EOS) for volatiles for which we use an EOS for water (Haldemann2020) which covers all phases. For simplicity, we treat for the evolution phase all volatile species accreted as ice (i.e. H2O, CO, CO2, CH3OH, CH4, NH3, N2, and H2S, see Marboeuf2014b; Marboeuf2014) as water ice and will use ”water” and ”volatiles” as synonyms in this work.
@@ 2.2.1 Planetary structure and composition
In all cases, we start at the disk dissipation time with initial conditions from the formation part of the models (Sect. 2.1). To improve the consistency, we recover the luminosities of the planets at analytically and rerun the evolution part from this time onward. A difference to Burn2024 is that we also use the hydrogen and helium (H/He) mass from as starting condition instead of the one after of evolution. This has an effect on simulated planets for which an impact occurred after . By starting with , we assume that no H/He is lost for the more massive planet in a collision while all H/He is lost for the impacting planet. The mass of other components, that is iron, rocky materials, and volatile ices, is however assumed to perfectly merge. Future works will re-introduce a more realistic impact stripping with the possibility of mass loss of all species. Therefore, in the scenarios where we assume that volatile species mix with H/He (Mixed, Fractionation, see below) we obtain a total heavy element content (by mass) in the envelope of
smaller or equal to that in Burn2024.
Further common elements in all models are the presence of an iron core with a silicate MgSiO3 (perovskite) mantle both materials described with a modified polytropic equation of state of Seager2007.
In all model cases, the uppermost – third or fourth – layer is what we call the envelope and is numerically resolved in one dimension. Similar to the formation stage (described in Emsenhuber2020a), the following internal structure equations (e.g. Kippenhahn2012) in hydrostatic equilibrium are solved:
@@
is smaller than the radiative gradient. The latter depends on the local optical depth . If , we use the atmosphere solution of Guillot2010 which considers an outgoing infrared flux and a well separated in-going flux in the visible wavelength regime in a in-parts gray atmosphere under the Eddington approximation (see also Jin2014). An important parameter in that model is the ratio of the opacities for the incoming (visible) and outgoing radiation (thermal). Here, we use linear interpolation within the values tabulated as a function of equilibrium temperature by Jin2014 in their Table 2. If the equilibrium temperature is outside the tabulated interval (260 to 2777 K), we extrapolate linearly below 260 K and quadratically in above 2777 K. We emphasize that the opacity ratios were obtained from detailed radiation transfer models assuming Solar composition envelopes. If sub-Neptunes are water rich, we would expect stronger greenhouse effect due to high infrared opacities of water. In this work, we neglect this effect, which allows us to isolate the effects of photoevaporation, which are expected to dominate for the short-orbit planets considered here.
For the explicit dependency on optical depth in the solution of Guilera2010, as well as for the radiative diffusion below, we use metallicity dependent opacities tabulated by Freedman2014 but note that there is a maximum tabulated metallicity of [M/H] =1.7, that is, it is not strictly applicable to . Both the opacity and their incoming to outgoing ratio should be generalized to differing compositions in future studies. In case of large optical depths, radiative diffusion is a well justified approximation, thus
@@
Following Mordasini2012c, we assume that the luminosity is constant throughout the structure which reduces the number of equations and is justified as only the upper, radiative regime depends directly on the luminosity. The system of equations is closed by the EOS (see below). A shooting method is used to solve the boundary value problem which arises due to core size, mass, and temperature being given as lower boundary while external pressure and density are upper boundaries to the envelope structure. A dependency on the stellar type enters directly into the upper boundary temperature of the planetary structures in the form of the irradiation temperature , where the stellar temperature and radius are taken from the tracks of Baraffe2015.
Assuming that a hydrostatic equilibrium is established on short timescales, the structure of the planet is recalculated after each step forward in time. From the obtained pressure and temperature profile, the corresponding change in energy content (gravitational and internal) can then be estimated. This estimate can be used as the luminosity for the next timestep and leads to a method which was found to result in a consistent treatment of a cooling and contracting planet (Mordasini2012c). In addition to the energy budget from solid and gas accretion, we consider radiogenic heating from the decay of 40K, 232Th, and 238U (Mordasini2012b) as an additional energy source. We further add an empirically determined source of energy causing the observed bloating of close-in planets following Sarkis2021.
@@
As for the solid part of the planetary structure, we require an equation of state for the envelope matter. For the mixed model, the envelope is assumed to consist of a perfect mixture of water, hydrogen, and helium. Therefore, the envelope metallicity is set to . While perfect mixing at all conditions is an extreme assumption, there are clear indications that H/He and water are well miscible at the high temperatures within the runaway greenhouse limit (Vazan2022; Pierrehumbert2023), which is a safe assumption for the observed population of small transiting exoplanets. The equation of state for water follows Haldemann2020 while that for H/He follows Chabrier2019. To combine them, we assume that Amagat’s law of additive volumes holds.
@@
In the Layered model, water and H/He are assumed to not mix at all to explore the other fringe case. We consider this model interesting from a theoretical perspective but physically less likely. Further, it resembles the classical onion-like structure assumption often used in the field. In this case, the resolved planetary structure is split into two part, an upper, pure H/He layer on top of the layer including all volatiles, treated as pure water. The uppermost layer follows the general structure outlined above. The pure water layer is solved assuming that the structure is adiabatic (i.e. solving Eqs. 2, 3, 4 but using the adiabatic gradient – Eq. 5 – in any case in Eq. 4) and employs the AQUA equation of state to close the system (Haldemann2020). It uses a fixed 2000 layers and obtains the solution for a given external pressure and temperature, obtained from the previous iteration of the H/He layer, via bisection. The interplay of the two structures is handled by fivefold global iteration per timestep, which is sufficient in phases of slow evolution. Testing different iteration procedures only gave slight differences in the timing (by a few years) of complete loss of H/He, thus we consider this approach acceptable.
After the loss of H/He (i.e. falling below a threshold mass of ), we switch for the volatiles to a pure water envelope with the full set of the general treatment but without H/He, thus using pure AQUA EOS. This transition implies for hot planets that we first force the water to be adiabatic, motivated by the pressure of the H/He on top of it, and then transform it to a typically supercritical vapor envelope which is partially radiative. We note that for planets which do not contain volatile ices or H/He after formation the Layered and Mixed models are equivalent.
@@
The structure and initialization of the Fractionation model is equal to the Mixed model explained above. The difference lies in loosening the assumption that the gas is perfectly mixed also in the photoevaporative wind. Instead, an analytic prescription (Hunten1987; Zahnle1986; Zahnle1990) detailed in Appendix A (optimistic estimate)222We use the analytic estimates from Appendix A and the expressions for there instead of the output from the tables of Johnstone2020 (Sect. 2.2.2) for its broader applicability. is used to determine the metallicity in the photoevaporative mass loss . Consequently, this also implies that can evolve, that is, increase, over time.
@@
A novelty in evolutionary calculations is the here presented Water Sequestration model. Instead of mixing the accreted volatile species (treated as water) mass only within the envelope, we maximize the amount of water sequestered in the deep interior by assuming the presence of a molten magma ocean during the whole evolution of the planet. In that case, some amount of volatiles can dissolve into the magma and can even dissolve into the metallic core based on the calculations in Luo2024 (see also Dorn2021). The envelope volatile mass at the start of our long-term evolution phase is then only and the envelope metallicity . The rationale of the model is to provide an upper limit for the mass of volatile species, treated as water, which could be removed from the envelope mass budget. This has implications for the evolution of the planetary envelope. To provide this upper limit of the effect, we assume a completely molten mantle (see discussion in Sect. 4.3) which can incorporate orders of magnitude more volatile mass compared to a solid mantle (Shah2021).
The partitioning of water between the mantle and surface water layers is modeled using a modified version of Henry’s law which relates the water mass fraction in the magma to the pressure at the bottom of the gaseous envelope (see Dorn2021 for details).
The parameters and were fitted to experimental data by Dorn2021. Partitioning from the magma ocean to the metallic core (to determine ) follows Luo2024 who assume iron silicate equilibration at mid-mantle pressures. Their partitioning coefficients depend on the concentration of water in the molten silicates as well as mid-mantle pressures but only marginally on the temperature at a given pressure.
For the coupling to the evolutionary cooling and photoevaporation model, tabulated results from these partitioning and solubility laws and further using the interior structure model detailed in Appendix B were used. The planetary structure for which a solution is found when creating the tables includes core, mantle, and water atmosphere. Thus, a consistent planet structure with water partitioning to core and mantle can be found via iteration for a given total mass and water mass fraction. We note that we did not include an explicit temperature dependency on the water partitioning leading to a weak effect of external temperature via the pressure dependencies. Therefore, the tables were using an exterior temperature of 3000 K for all models and varying total water mass budgets (from 1% to 30%) and planetary masses (0.5 to 30 M⊕) were considered. For the equilibration of water between silicates and iron, we use mid-mantle pressures and the iron mass fraction of the dry core was assumed to be 0.244 consistent with the outcome of the equilibrium condensation model used in our formation model (Thiabaud2014). To construct the tables, the presence of H/He, which would increase pressures and thus solubilities, was excluded for simplicity but supported by the fact that most planets in the considered size regime contain little H/He in mass.
@@
During time evolution, the total mass of the planet and the atmospheric pressure can decrease due to photoevaporative mass loss. Therefore, the balance between sequestered water in the magma ocean and envelope (Eq. 7) is re-evaluated after each timestep and is re-determined using the same tables as initially. The water partitioning between envelope and magma ocean is therefore dynamically adapted to the new conditions which leads to outgassing of water from the magma and a gradual enrichment of the envelope with water vapor until is reached. In contrast, we assume that the volatiles trapped in the iron core remain constant over time (similar to outgassed scenarios in Rogers2024a). The same tabulated data can be used because the water reservoir in the magma ocean is not affected by the amount of water in the core but by exterior conditions.
For simplicity, the Water Sequestration model calculates the average density of the volatile-enriched metallic core and magma ocean using the equation of state of Seager2007 assuming the same separation into iron core, perovskite mantle, and water ice layer as in the Layered model. This implies that for the radius and bulk density of the interior, we do not assume that sequestration to the mantle and core occurred. Instead, the sequestered water (of mass ) is treated as a high-pressure ice layer on top of the rocky material. A comparison to densities calculated consistently with the structure assumption of the Water Sequestration model (Fig. 1) is presented in Appendix B.
@@ 2.2.2 Photoevaporative mass loss
To model the escape of particles from the upper atmosphere due to high-energy radiation, we use a similar prescription for photoevaporation as in Affolter2023 and Burn2024. We assume that the escape is in the thermally-driven hydrodynamic regime (see Gronoff2020, for a recent review). The total mass loss is a combination of the loss of H/He (Kubyshkina2021) and that of volatile species treated as water (Johnstone2020). Given the , differing between models (see Sect. 2.2.1), it is given by
where we use the tabulated values of Kubyshkina2021 for . The results are based on the chemical-hydrodynamic models of Kubyshkina2018, assume a 15% heating efficiency of the high-energy radiation, and depend on stellar mass, X-ray and extreme ultraviolet (EUV) fluxes , planetary masses and radii, as well as equilibrium temperature. To transition from these models applicable for pure H/He planets to water-dominated cases, the results of Johnstone2020 are used to determine in Eq. (8) by adjusting the efficiency parameter in the energy-limited mass loss equation
where is the received flux at the planet location in either X-ray (dominating early) or EUV wavelengths, is the radius of the layer where an optical depth of 2/3 is reached from the inside-out, and is the base of the ionization layer. It is found by searching within the resolved planetary structure for the location where an optical depth of one if reached for UV photons (Murray-Clay2009). The factor , with being the ratio of the planets Roche limit to its radius. As in Burn2024, the time evolution of follows (Mcdonald2019) and we linearly interpolate or extrapolate to lower stellar masses if required. In contrast to Burn2024, we reduced the rates of Johnstone2020 to account for an additional cooling effect fitted following Yoshida2022 as , where is the unmodified rate fit to the results of Johnstone2020. We note that the work by Yoshida2022 was conducted using the XUV spectrum of an M8 star (motivated by TRAPPIST-1), which differs significantly from a Solar-type star and is therefore most appropriate at the low stellar mass end.
@@ 2.3 Observational data and sampling
To compare simulations and observations, we use the collection of well characterized exoplanets from Parc2024 with the goal of comparing mass-radius relations. For our purposes, we restrict ourselves to planets within 30 days orbital periods and smaller than 7 R⊕. Furthermore we require precisely determined masses and radii with relative errors below 25% and 8% respectively. This results in a set of 201 observed planets. As a second exploration, we will relate the Solar-mass population of simulated exoplanet radii to the data from the California-Kepler survey (Fulton2018). For this second set of data, no precise masses are available but the application of the bias from Kepler can be done in a rigorous fashion using the KOBE package Mishra2021.
@@
To allow for a tentative, quantitative comparison also for the sample from Parc2024, we apply an estimate for the observational bias and re-sample the synthetic population that is an outcome of the model described above such that the distribution of stellar masses and the right number of observed stars is reproduced. The observational data originates from heterogeneous sources and a human intervention bias is introduced when follow-up characterization observations are scheduled. Nevertheless, a significant fraction of the planets were discovered by TESS (Ricker2014) which is a transit survey. Therefore, we calculate the geometrical transit probability (Petigura2018)
@@
and an estimate for the probability that a transiting planet is also detected (based on the Kepler mission, as detailed by Fulton2017; Petigura2018)
@@
where is the standard cumulative distribution function, , and . Here, we use the expected scaling of signal to noise with observation time (Petigura2018), estimate the mission duration ratio of Kepler to TESS , and increase by a factor of two to account for the heterogeneous sample as well as the variability in TESS observation duration per start (depends on the sector). We visually verify the estimated detection probabilities by comparing them to the results of the detailed analysis of TESS planet detection yields by Kunimoto2022.
@@
Most of the models (except the Layered one) rely on the assumption that water mixes with H/He. This is established for planets which are interior to the runaway greenhouse limit where water cannot condense due to too high temperatures. Therefore, we apply a cut at the runaway greenhouse limit which affects planets around different stars differently: The runaway greenhouse limit lies at close orbits similar to the detection thresholds for late M dwarfs only. For the estimate we use either the observed luminosities for the observational data or for synthetic stars, we use stellar luminosities at from the tracks by Baraffe2015. The incident flux of shortwave radiation on the planetary surface is on average , where is the planets’ albedo. Assuming that the planet is in thermal equilibrium and for a limiting case assuming no internal heat content, we can equate this to a critical outgoing long-wave radiation limit of from Boukrouche2021. The resulting critical distance
@@
is then a conservative upper limit within which our assumption that the water evaporates and mixes with other constituents holds. To remain conservative, we use a relatively high albedo of motivated by the results of Kopparapu2014 for high-pressure envelopes.
@@
To generate a synthetic sample of planets we sample systems from the original synthetic data weighted by stellar mass chosen such that the stellar mass distribution of the synthetic data matches the observed one (equivalent to Schlecker2022). We then proceed to draw “observed” planets by repeatedly (1000 times a set of 201) sampling planets weighted by their individual thus allowing for an individual, modeled planet to be present several times in the synthetic comparison sample. We note that this approach ignores possible correlations of the detection probability within a planetary system and we used discrete stellar mass bins (0.1, 0.3, 0.5, 0.7, 1.0, and ) when the synthetic data was created (Burn2021).
@@
Since the observed planets, to which we want to compare our models, are also characterized by radial velocity measurements, we further require for planets to be included in the synthetic comparison data that the semi-amplitude (e.g. Cumming1999, their Eq. 1) is larger than .